OpenStudy (anonymous):

Solve the systems using substitution 6x + 2y = 0 4x + y = -1

OpenStudy (angela210793):

4x + y = -1 --->y=-1-4x 6x+2(-1-4x)=0 6x-2-8x=0 -2x=2 x=-1 y=-1-4*(-1)=-1+4=3 solution is (-1;3)

OpenStudy (anonymous):

ty!